what id the formula of power

what id the formula of power
Answer #1

There are multiple ways to define power. Based on these definitions, we have different formula of power.

Note that despite these different formulae, power still has the dimensions of energy per second.

Work-Energy Definition

Power is defined here as the work done in unit time, or the energy transferred per unit unit time.

P=\frac{W}{t}

Force-Velocity Definition

Here, Power is defined as rate of the force applied, as the object travels.

P=Fv

We can also derive it from the previous definition, by realising that W=Fx , and v=\frac{x}{t} .

\\ P=\frac{W}{t} \\ P=\frac{Fx}{t}=Fv

Electrical Energy Definition

Here, power is the amount of charge transferred in unit time across a potential difference.

The amount of charge transferred in a unit time is the current, and using this, we get the next formula.

P=VI